# Adaptive Budgeted Multi-Armed Bandits for IoT with Dynamic Resource Constraints

### v1

@@ I-A Related Work
Recent advances have explored constrained multi-armed bandits (CMAB), where the agent must optimize rewards while satisfying fixed constraints. Pioneering works such as those by Badanidiyuru et al. [4] introduced algorithms like Conservative Bandits. Similarly, Ding et al. [5] proposed primal-dual methods for constrained bandits. However, these methods do not address dynamically evolving constraints.
Another line of research investigates safe exploration techniques, such as Safe-UCB by Sui et al. [6] and more recent approaches leveraging shielding in constrained MDPs [7]. These approaches typically assume static constraints or conservative assumptions that limit reward optimization.
Work on multi-objective bandits such as Drugan and Nowé [8] extends bandit models to multiple competing objectives, but usually without asymmetric roles between reward and constraint. They do not differentiate between the primary objective (e.g., throughput maximization) and secondary operational constraints (e.g., energy consumption). More recent work, like Berkenkamp and Krause [9], integrates constrained optimization for safe online learning.
Reinforcement learning acts as a promising approach for such dynamically changing and constrained 6G IoT environments, owing to its ability to learn optimal policies while interacting with the environment [10]. Recent studies (e.g., Li et al. [11] and Wang and Sun [12]) highlight constrained learning for energy-critical and 6G IoT environments, further underscoring the need for adaptive constraint handling mechanisms. However, these models often assume perfect or predictable knowledge of environmental constraints and lack the flexibility required to adapt to stochastic and time-varying constraints and budgets.
While Neely and Yu’s OCO algorithm [13] develops a virtual-queue method for full-information convex optimization with time-varying constraints under i.i.d. assumptions, our work addresses the partial-information MAB setting with stochastic reward and cost feedback. Cao and Liu [14, 15] study online convex optimization with time-varying constraints, addressing both full information and bandit feedback settings. Their algorithms achieve sublinear regret and constraint violation, assuming sublinear drift of the comparator sequence. We introduce a decaying violation budget and propose a Budgeted UCB algorithm that allows controlled exploration-phase violations yet enforces vanishing constraint breaches over time, obtaining novel convergence guarantees in the dynamic-IoT bandit context.
@@ IV Theoretical Results
With probability at least $1-1/T$, Budgeted UCB satisfies
@@
The exact regret and violation bounds for the exploration phase of Budgeted UCB are derived from the classical analysis of the Upper Confidence Bound (UCB) algorithm found in the foundational work of Auer et al. (2002) [3]. Specifically, with probability $1-1/T$, the bound on regret in the exploration phase is given explicitly as $R_{exp}(T)$:
@@
Note that the violation budget, $\delta_{t}$, reduces to 00 after a constant $T_{bud}$ exploration period.
@@
Part 3 - Total Regret Bound. With probability at least $1-1/T$,
@@ V-A Simulation Setup and Baselines
We consider a scenario of an IoT device, a battery operated transmitter which sends data over a wireless channel to a fixed receiver located $d=10$m away. Our aim is to maximize the cumulative throughput $\sum_{t=1}^{T}r_{t}$ over a horizon of $T=2000$ time steps, while keeping the empirical energy‐constraint violation rate $v_{t}$ below a linearly decaying budget $\delta_{t}$:
@@
where $\delta_{0}=0.5$. At each round an energy constraint $C_{t}$ is imposed, as shown in Figure 1. We conduct two experiments: in Experiment 1, $C_{t}$ is drawn uniformly in $[P_{\min},P_{\max}]$, and in Experiment 2 drifting linearly down from $P_{\max}$ to $P_{\min}$ then back up. Our Budgeted UCB algorithm uses $\delta_{t}$ to decide when to explore purely for throughput versus when to filter arms by their cost‐UCB to avoid violations.
The wireless communication channel has bandwidth $B=1$MHz and experiences path‐loss $g=d^{-\alpha}$ with exponent $\alpha=3$, plus additive white Gaussian noise of spectral density $N_{0}=10^{-9}\,\mathrm{W/Hz}$. To adapt to changing energy constraints, the transmitter may choose at each time $t$ one of $K$ discrete power levels $P(a)\;\in\;\{P_{\min},\,P_{\min}+\Delta P,\;\dots,\;P_{\max}\},$ where $P_{\min}=0.1$ W, $P_{\max}=1.0$ W and $\Delta P=(P_{\max}-P_{\min})/(K-1)$. Transmitting at power $P(a_{t})$ yields an instantaneous data rate (throughput) and energy cost
@@
Bandwidth $B$ $1$ MHz Noise Density $N_{0}$ $10^{-9}$ W/Hz
Path-loss exponent $\alpha$ $3$ Number of Arms $11$

### v2

@@ I-A Related Work
Recent advances have explored constrained multi-armed bandits (CMAB), where the agent must optimize rewards while satisfying fixed constraints. Pioneering works such as Badanidiyuru et al. [4] proposed two algorithms for the Bandits with Knapsacks problem - BalancedExploration and PrimalDualBwk - that achieve near-optimal regret under fixed, known resource constraints. However, their framework assumes static budgets and does not address scenarios with dynamically evolving constraints.
Another line of research investigates safe exploration techniques, such as Safe-UCB by Sui et al. [5] and Moradipari et al. [6]. These approaches typically assume static constraints or conservative assumptions that limit reward optimization.
Multi-objective bandits approaches, such as Drugan and Nowé [7] extend bandit models to handle multiple competing reward dimensions, optimizing for trade-offs via Pareto-based or scalarization methods. These methods treat all objectives symmetrically and do not differentiate between a primary performance objective (e.g., throughput) and secondary operational constraints (e.g., energy use). In contrast, a more recent work by Wang et al. [8] integrates constrained optimization for safe online learning.
Recent studies-such as Chen and Giannakis [9] for energy-critical IoT systems and Liu and Fang [10] for 6G IoT task scheduling—demonstrate the value of incorporating constraints into online learning and decision-making. While these models address important operational limits, they typically assume perfect or predictable knowledge of constraints and thus lack the flexibility needed to handle stochastic, time-varying budgets and requirements.
While Neely and Yu’s OCO algorithm [11] develops a virtual-queue method for full-information convex optimization with time-varying constraints under i.i.d. assumptions, our work addresses the partial-information MAB setting with stochastic reward and cost feedback. Cao and Liu [12, 13] study online convex optimization with time-varying constraints, addressing both full information and bandit feedback settings. Their algorithms achieve sublinear regret and constraint violation, assuming sublinear drift of the comparator sequence.
Reinforcement learning is a promising approach for these dynamically changing and constrained 6G IoT environments, due to its ability to learn optimal policies while interacting with the environment [14]. We introduce a decaying violation budget and propose a Budgeted UCB algorithm that allows controlled exploration-phase violations yet enforces vanishing constraint breaches over time, obtaining novel convergence guarantees in the dynamic-IoT bandit context.
@@ IV Theoretical Results
With probability at least 11/T1-1/T, Budgeted UCB satisfies
@@
The exact regret and violation bounds for the exploration phase of Budgeted UCB are derived from the classical analysis of the Upper Confidence Bound (UCB) algorithm found in the foundational work of Auer et al. (2002) [3]. Specifically, with probability 11/T1-1/T, the bound on regret in the exploration phase is given explicitly as $R_{exp}(T)$:
@@
Note that the violation budget, $\delta_{t}$, reduces to 0 after a constant $T_{bud}$ exploration period.
@@
Part 3 - Total Regret Bound. With probability at least 11/T1-1/T,
@@ V-A Simulation Setup and Baselines
We consider a scenario of an IoT device, a battery operated transmitter which sends data over a wireless channel to a fixed receiver located d=10d=10m away. Our aim is to maximize the cumulative throughput $\sum_{t=1}^{T}r_{t}$ over a horizon of $T=2000$ time steps, while keeping the empirical energy‐constraint violation rate $v_{t}$ below a linearly decaying budget $\delta_{t}$:
@@
where $\delta_{0}=0.5$. At each round an energy constraint $C_{t}$ is imposed, as shown in Figure 1. We conduct two experiments: in Experiment 1, $C_{t}$ is drawn uniformly in $P_{\min},P_{\max}]$, and in Experiment 2 drifting linearly down from $P_{\max}$ to $P_{\min}$ then back up. Our Budgeted UCB algorithm uses $\delta_{t}$ to decide when to explore purely for throughput versus when to filter arms by their cost‐UCB to avoid violations.
The wireless communication channel has bandwidth B=1B=1MHz and experiences path‐loss $g=d^{-\alpha}$ with exponent $\alpha=3$, plus additive white Gaussian noise of spectral density $N_{0}=10^{-9}\,\mathrm{W/Hz}$. To adapt to changing energy constraints, the transmitter may choose at each time $t$ one of $K$ discrete power levels $P(a)\;\in\;\{P_{\min},\,P_{\min}+\Delta P,\;\dots,\;P_{\max}\},$ where $P_{\min}=0.1$ W, $P_{\max}=1.0$ W and $\Delta P=(P_{\max}-P_{\min})/(K-1)$. Transmitting at power $P(a_{t})$ yields an instantaneous data rate (throughput) and energy cost
@@
Bandwidth $B$ 11 MHz Noise Density $N_{0}$ $10^{-9}$ W/Hz
Path-loss exponent $\alpha$ 33 Number of Arms $11$